Circular No. 8738

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMETS C/2006 K7–K21, C/2006 L3–L5 (SOHO) Additional Kreutz sungrazing comets have been found on SOHO web-

site images (cf. IAUC 8727, 8737; LC = L. Cane). C/2006 K7 is described by K. Battams as being slightly diffuse and small, reaching mag 8.1 on May 18.313 UT at 6.1R_{⊙}. C/2006 K8 was faint, quite diffuse, and too faint for photometry. C/2006 K9 was extremely faint and diffuse. C/2006 K10 was very faint, diffuse, and slightly elongated. C/2006 K11 appeared small and stellar in C3 images, while in C2 images it reached mag ∼ 6.5 and had a tail that reached a length of 113′′ on May 23.254. C/2006 K12 was small, diffuse, and very faint. C/2006 K13 and K14 were a fairly close pair — both being very small and faint — the former (the leading fragment) being diffuse and the brighter of the two, and the latter being slightly diffuse. C/2006 K15, K16, K18, K19, and C/2006 L5 were all faint and slightly diffuse; C/2006 K15 and K19 were described as tiny, the other two as small. C/2006 K17 and K20 were diffuse, the former extremely faint and the latter very small. C/2006 K15–20 all peaked in the magnitude range 8–9. C/2006 K21 was small and stellar in appearance, peaking at mag ∼ 4.5 in C3 images at ∼ 13R_{⊙}; there was a vague hint of a tail in C2 images. C/2006 L3 was very faint and diffuse. C/2006 L4 was tiny and slightly diffuse.

Comet 2006 UT α_{2000} δ_{2000} Inst. F MPEC
C/2006 K7 May 18.314 3^{h}42.^{m}4 +18^{o}00′ C2  LC 2006-M28
C/2006 K8  18.898 3 45.2 +18 09 C2  BZ 2006-M28
C/2006 K9  19.913 3 47.6 +18 13 C2  RM 2006-M28
C/2006 K10  20.564 3 51.0 +18 26 C2  WX 2006-M30
C/2006 K11  22.513 4 03.3 +17 20 C3/2 HS 2006-M30
C/2006 K12  22.754 3 59.4 +18 50 C2  HS 2006-M30
C/2006 K13  24.421 4 05.8 +19 08 C2  GS 2006-M30
C/2006 K14  24.438 4 05.8 +19 09 C2  RM 2006-M30
C/2006 K15  25.004 4 07.7 +19 12 C2  TC 2006-M64
C/2006 K16  25.913 4 11.4 +19 22 C2  RM 2006-M64
C/2006 K17  25.938 4 11.5 +19 21 C2  RM 2006-M64
C/2006 K18  26.413 4 11.9 +19 24 C2  GS 2006-M64
C/2006 K19  29.300 4 23.9 +19 51 C2  HS 2006-M64
C/2006 K20  29.408 4 24.7 +19 54 C2  TC 2006-M64
C/2006 K21  29.554 4 32.8 +17 18 C3/2 VB 2006-O20
C/2006 L3  June 1.908 4 38.1 +20 21 C2  GS 2006-O20
C/2006 L4  4.592 4 48.2 +20 39 C2  WX 2006-O20
C/2006 L5  8.058 5 00.7 +20 56 C2  HS 2006-O20
2006 August 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green